integrate the following if possible

e3x?-8**l(3x – 4)dx
Answer #1

Given, , 237-8x+1 (3x-4) 64 X+1 Substitute, > 3x-8x11 st d(3x²-8x+1) = dt (64-8)dx=dt - (3x-4) dx = dt so, integra becomes; 2